OpenStudy (anonymous):

What is the midpoint of the line segment with endpoints (5, -2) and (1, -6)?Select one of the options below as your answer: A. (2, -2) B. (3, -4) C. (3, -6) D. (4, -3)

OpenStudy (anonymous):

(3,-4) b.

OpenStudy (lalaly):

( 5+1 / 2)= 3 so x= 3 -2-6 / 2= -4 so y= -4 (3,-4)